OpenStudy (snitchseeker1496):

The distance between the centers of the three circles, which are mutually tangent with each other, are 10,12 and 14 units, Find the area of the largest circle.

OpenStudy (snitchseeker1496):

@jim_thompson5910 can you help me ? :)

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

This is a tough problem and I had to do a bit of searching. I found this though https://www.ics.uci.edu/~eppstein/junkyard/tangencies/three-circles.html which helps construct the three circles This is what I get when I follow the steps in that article. See attached. Again I used geogbera to help me out

jimthompson5910 (jim_thompson5910):

As for how to do this without using a visual aid like this, I'm not sure. I have a feeling it will rely on using coordinate geometry somehow

OpenStudy (zarkon):

it can be solved with a simple system of equations

jimthompson5910 (jim_thompson5910):

@Zarkon do you mean this? x+y = 14 y+z = 10 x+z = 12 where x,y,z are the radii of the circles. I've thought about it but I overlooked it because it seemed too easy, but idk

OpenStudy (zarkon):

yes

jimthompson5910 (jim_thompson5910):

ok then go with that method @SnitchSeeker1496

jimthompson5910 (jim_thompson5910):

x+y = 14 y+z = 10 solve the second equation for y y = 10-z then use substitution x+10-z = 14 x-z = 4 ------------------------------- So we now have these two equations x-z = 4 x+z = 12 use elimination to solve

OpenStudy (snitchseeker1496):

x =16, y= -2, z = 12 ?

OpenStudy (snitchseeker1496):

@jim_thompson5910 after getting the the values of x, y and z, what should I do next?

jimthompson5910 (jim_thompson5910):

`x =16, y= -2, z = 12 ?` is incorrect unfortunately

OpenStudy (snitchseeker1496):

Oh wait, x is for the largest circle right? so the answer is 64 pi for the area?

jimthompson5910 (jim_thompson5910):

it sounds like you got 8 to be the largest radius which is correct, so yes, the largest circle has area 64pi

OpenStudy (snitchseeker1496):

i forgot to divide it by 2 sorry :))

jimthompson5910 (jim_thompson5910):

btw you should find that... x = 8 y = 6 z = 4

OpenStudy (snitchseeker1496):

x = 8, y = 6 , z =4 right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (snitchseeker1496):

Thank you :)

jimthompson5910 (jim_thompson5910):

no problem